OpenStudy (anonymous):

Least to greatest- 1/2 , 4/3 , 8/8 , 6/5 ,7/4

OpenStudy (anonymous):

let's reduce 8/8 to 1: 1/2, 4/3, 1, 6/5, 7/4 Now what's the lowest common denominator we can find? I think it's 60 (3 * 4 * 5): 1/2 = 30/60 4/3 = 80/60 1 = 60/60 6/5 = 72/60 7/4 = 105/60 Now order from least to greatest.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Lets see

OpenStudy (anonymous):

tell me if im right or wrong ok guys?

OpenStudy (anonymous):

better hurry, my battery is dying. :-)

OpenStudy (anonymous):

ummm can u just tell me the answer i got to go soon please i got school :(

OpenStudy (anonymous):

order the fractions from least to greatest by numerator (after they were converted to /60).

OpenStudy (anonymous):

the least is 1/2 (30/60)...

OpenStudy (anonymous):

1/2 , 8/8 , 4/3 , 6/5 , 7/4 ?

OpenStudy (anonymous):

close... is 80 less than 72?

OpenStudy (anonymous):

(you swapped two of them)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so first one is 1/2 i put?

OpenStudy (anonymous):

the answer is 1/2 , 8/8 , 6/5 , 4/3 , 7/4

OpenStudy (anonymous):

in order of numerator 30, 60, 72, 80, 105

OpenStudy (anonymous):

thank u sir your a good help

OpenStudy (anonymous):

umm before u go can u tell me the rule ? it says write a rule for ordering any fractions?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

my battery actually died. and I went to sleep. the rule is: convert all the fractions to have a common denominator, then order by numerator.